# Integral cohomology rings of weighted Grassmann orbifolds and rigidity properties

### 11 Mar 2024 v2

@@ 3. Classification of weighted Grassmann orbifolds upto certain homeomorphism
Two weighted projective spaces $\mathbb{W}P(b_{0},\dots,b_{m})$ and $\mathbb{W}P(c_{0},\dots,c_{m})$ are homeomorphic iff $(b_{0},\dots,b_{m})$ and $(c_{0},\dots,c_{m})$ have same normalized form upto an order.
Remark 3.8.
Let ${\bf b}=(b_{0},b_{1},\dots,b_{m})$ and ${\bf c}=(c_{0},c_{1},\dots,c_{m})$ be two Plücker weight vectors such that $\mbox{Gr}_{\bf b}(k,n)$ and $\mbox{Gr}_{\bf c}(k,n)$ be homeomorphic. If we can extend this homeomorphism to get a homeomorphism between $\mathbb{W}P(b_{0},\dots,b_{m})$ and $\mathbb{W}P(c_{0},\dots,c_{m})$ then the normalized forms of $(b_{0},\dots,b_{m})$ and $(c_{0},\dots,c_{m})$ are the same up to an order.
In the remaining part of this section we study the rigidity problems of $\mbox{Gr}_{\bf b}(k,n)$ for $k\geq 2$. Define
@@
Then cardinality of $Ad(\lambda^{i})$ is $k(n-k)$. Corresponding to $Ad(\lambda^{i})$ we define
${\bf b}_{(i)}:=\{{b_{j}}:j\in Ad(\lambda^{i})\}\text{ and }\overline{\bf b}_{(i)}:={\bf b}_{(i)}\cup\{b_{i}\}.$
Consider $[{e}_{\lambda^{i}}]\in\mbox{Gr}_{\bf b}(k,n)$, where $e_{\lambda^{i}}=(0,\dots,0,\underset{i^{th}}{1},0,\dots,0)\in Pl(k,n)$. We call the elements $[e_{\lambda^{i}}]$ by the coordinate elements, for $i=0,1,\dots,m$. Let
(3.4) $U_{i}=\{[z_{0},\dots,z_{m}]\in\mbox{Gr}_{\bf b}(k,n)~{}|~{}z_{i}\neq 0\}.$
From Section 2, it follows that $U_{i}\cong\mathbb{C}^{k(n-k)}/G(b_{i})$. Note that $U_{i}$ is an open cone over $L^{\prime}(b_{i};{\bf b}_{(i)})$. Let $t:=k(n-k)$. Using [13, Theorem 2] we have
$H^{q}(L^{\prime}(b_{i};{\bf{b}}_{(i)});\mathbb{Z})=\begin{cases}\mathbb{Z}&\quad\text{for}~{}q=0,2t-1\\ \mathbb{Z}_{\mu}&\quad\mu={l_{e}{\overline{\bf{b}}_{(i)}}}/{l_{e}{{\bf{b}}_{(i)}}}~{}\text{if}~{}q=2e~{}\text{for}~{}1\leq e\leq t-1\\ \{0\}&\quad\text{ otherwise}.\end{cases}$
@@
(3.5) $l_{e}{\bf{b}}_{(i)}=\prod_{p~{}\text{prime}}\prod_{i=1}^{e}p^{r_{t-i+1}}~{}\text{for}~{}1\leq e\leq t-1.$
@@
If ${\bf b}$ is a primitive Plücker weight vector then $\overline{\bf{b}}_{(i)}$’s are primitive vectors for all $i=0,1,\dots,m$.
@@
Fix one $i\in\{0,1,\dots,m\}$. If $b_{j}\notin\overline{\bf b}_{(i)}$ then $|\lambda^{j}\cap\lambda^{i}|=k-u\leq k-2$. Without loss of generality assume that $\lambda^{i}=(\lambda^{i}_{1},\dots,\lambda^{i}_{u},\lambda^{i}_{u+1}\dots,\lambda^{i}_{k})$ and $\lambda^{j}=(\lambda^{j}_{1},\dots,\lambda^{j}_{u},\lambda^{i}_{u+1}\dots,\lambda^{i}_{k})$, where $\lambda^{i}_{1}\neq\lambda^{j}_{1},$ $\dots$,… , $\lambda^{i}_{u}\neq\lambda^{j}_{u}$. Consider Schubert symbols $\lambda^{s_{1}}=(\lambda^{j}_{1},\lambda^{i}_{2}\dots,\lambda^{i}_{u},\lambda^{i}_{u+1}\dots,\lambda^{i}_{k})$ $,\dots,\lambda^{s_{u}}=(\lambda^{i}_{1},\dots,\lambda^{i}_{u-1},\lambda^{j}_{u},\lambda^{i}_{u+1},\dots,\lambda^{i}_{k})$. Then $|\lambda^{i}\cap\lambda^{s_{v}}|=k-1$ and $b_{s_{v}}\in{\bf b}_{(i)}$ for all $v=1,2,\dots,u$. Note that
@@
If $d$ divides $b_{i}$ and every element of ${\bf b}_{(i)}$ then $d$ also divides $b_{j}$. Thus $\overline{\bf b}_{(i)}$ is not a primitive vector contradicts that ${\bf b}$ is a primitive Plücker weight vector. ∎
@@
If ${\bf b}$ is a primitive Plücker weight vector then
@@
Recall $U_{i}$ from (3.4). Using excision
@@
Now we need to prove
@@
If there exists a homeomorphism $f$ between two weighted Grassmann orbifolds ${\rm{Gr}}_{\bf b}(k,n)$ and ${\rm{Gr}}_{\bf c}(k,n)$ such that for all $i=0,1,\dots,m$, $f$ maps every coordinate element $[e_{\lambda^{i}}]$ to some coordinate element $[e_{\lambda^{j}}]$ then ${\bf b}$ and ${\bf c}$ are same up to a scalar multiplication and a permutation.
@@
We recall the (*$(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\bf b}(k,n)$ from (3.1) and note that the fixed point set of $\mbox{Gr}_{\bf b}(k,n)$ with respect to the (*$(\mathbb{C}^{*})^{n}$-action is $\{[e_{\lambda^{i}}]\}_{i=0}^{m}$.
@@
[…]
@@ 4. Torsion’s in the integral cohomology of weighted Grassmann orbifold
In this section, we prove that for a prime $p$ the integral cohomologies of weighted Grassmann orbifolds have no $p$-torsion under some assumptions. We prove that for any Plücker weight vector ${\bf b}\in(\mathbb{Z}_{\geq 1})^{6}$, $H_{i}(G_{\bf b}(k,n))$ is torsion free if $i\neq 3$. We discuss some non-trivial classes of weighted Grassmann orbifolds whose integral cohomologies have no torsions. In particular, the integral cohomology of a divisive weighted Grassmann orbifold is torsion-free and concentrated in even degrees.
@@
For a prime $p$, if there exists a Plücker permutation $\sigma$ on $\{0,\dots,m\}$ such that the for all $j\geq 3$, the $p$-th content of $b_{\sigma(j)}$ divides at least $d(\lambda^{j})-1$ many elements of the set $\{b_{\sigma(\ell)}:\lambda^{\ell}\in R(\lambda^{j})\}$ then the integral cohomology of ${\rm{Gr}}_{\bf b}(k,n)$ has no $p$-torsion.
@@
Consider the Plücker permutation $\sigma$ and recall the building sequence
$\sigma X^{0}\subset\sigma X^{1}\subset\dots\subset\sigma X^{m}=\mbox{Gr}_{\bf b}(k,n),$
induced from the $q$-CW complex structure as in Theorem 3.4, where $X^{i}$ is defined in (2.12). Note that $\sigma X^{i}$ is weighted projective space for all $0\leq i\leq 2$. Thus, the integral cohomology of $\sigma X^{i}$ has no torsion for all $0\leq i\leq 2$ by [13, Theorem 1]. We complete the proof by induction on $i$. Let $H^{*}(\sigma X^{i-1};\mathbb{Z})$ has no $p$-torsion. Consider the cofibration as in Proposition 2.12
$L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)})\hookrightarrow\sigma X^{i-1}\to\sigma X^{i},$
where ${\bf b}^{(i)}$ is defined in (2.13). This will induce the following exact sequence in cohomology
(4.1) $\to\widetilde{H}^{q-1}(L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)});\mathbb{Z})\to H^{q}(\sigma X^{i};\mathbb{Z})\to H^{q}(\sigma X^{i-1};\mathbb{Z})\to\widetilde{H}^{q}(L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)});\mathbb{Z}).$
Using [13, Theorem 2] we have
$H^{q}(L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)});\mathbb{Z})=\begin{cases}\mathbb{Z}&\quad\text{for}~{}q=0,2d(\lambda^{i})-1\\ \mathbb{Z}_{\mu}&\quad\mu={l_{e}{\sigma\overline{\bf{b}}^{(i)}}}/{l_{e}{\sigma{\bf{b}}^{(i)}}}~{}\text{if}~{}q=2e~{}\text{for}~{}1\leq e\leq d(\lambda^{i})-1\\ \{0\}&\quad\text{ otherwise},\end{cases}$
where ${\overline{\bf{b}}^{(i)}}:={\bf{b}}^{(i)}\cup\{b_{i}\}.$
For any prime $p$ the $p$-th content in $l_{e}{\sigma{\bf{b}}^{(i)}}$ is the product of $e$ many maximum $p$-th contents of the elements in set $\{b_{\sigma(\ell)}\colon\lambda^{\ell}\in R(\lambda^{i})\}$ as in (3.5). Thus the $p$-th content of $\mu=1$ if the $p$-th content of $b_{\sigma(i)}$ divides $e$ many element of the set $\{b_{\sigma(\ell)}:\lambda^{\ell}\in R(\lambda^{i})\}$. Now the maximum value of $e$ can be $d(\lambda^{i})-1$. Thus by our assumption $H^{*}(L^{\prime}(b_{\sigma(i)};\sigma{\bf{b}}^{(i)});\mathbb{Z})$ has no $p$-torsion. Thus $H^{*}(\sigma X^{i},\mathbb{Z})$ has no $p$-torsion using (4.1). Hence, by induction the integral cohomology of $\mbox{Gr}_{\bf b}(k,n)$ has no $p$-torsion. ∎
@@
Recall the subcomplexes $X^{i}$ of $\mbox{Gr}_{\bf{b}}(k,n)$ from (2.12) for $i=0,1,\dots,m$. Define
$\eta=\frac{\operatorname{lcm}\{b_{0},b_{1},b_{3}\}}{\operatorname{lcm}\{b_{0},b_{1}\}}\text{ and }\eta^{\prime}=\frac{\operatorname{lcm}\{b_{0},b_{1},b_{2}\}}{\operatorname{lcm}\{b_{0},b_{1}\}}.$
Lemma 4.3.
Proof.
Consider the cofibration
$L^{\prime}(b_{3};(b_{0},b_{1}))\to\mathbb{W}P(b_{0},b_{1},b_{2})\to X^{3}.$
This induces an exact sequence in homology by the following
(4.2) $\to\widetilde{H}_{i}(L^{\prime}(b_{3};(b_{0},b_{1})))\to H_{i}(\mathbb{W}P(b_{0},b_{1},b_{2}))\to H_{i}(X^{3})\to\widetilde{H}_{i-1}(L^{\prime}(b_{3};(b_{0},b_{1})))\to$
For $i=1$ in (4.2) we get
$0\to H_{1}(X^{3})\to\mathbb{Z}\to\mathbb{Z}\to.$
$H_{1}(X^{3})$ is either $\{0\}$ or a finite abelian group as $X^{3}$ has no cell of real dimension $1$. Also $H_{1}(X^{3})\to\mathbb{Z}$ is injective. Thus $H_{1}(X^{3})=\{0\}$.
For $i=3$ in (4.2) we get
$\to\mathbb{Z}\to 0\to H_{3}(X^{3})\to 0\to\mathbb{Z}.$
Thus $H_{3}(X^{3})=\{0\}$. For $i=4$ in (4.2) we get
$\to 0\to\mathbb{Z}\to H_{4}(X^{3})\to\mathbb{Z}\to 0.$
Thus $H_{4}(X^{3})=\mathbb{Z}\oplus\mathbb{Z}$. For $i\geq 5$ in (4.2) we get
$\to 0\to 0\to H_{i}(X^{3})\to 0\to.$
Now put $i=2$ in (4.2) we get
$\to 0\to\mathbb{Z}\to H_{2}(X^{3})\to\mathbb{Z}_{\eta}\to 0,$
$L^{\prime}(b_{2};(b_{0},b_{1}))\to\mathbb{W}P(b_{0},b_{1},b_{3})\to X^{3}.$
Considering the homology exact sequence corresponding to this cofibration, we can say $H_{2}(X^{3})$ has maximum $\eta^{\prime}$ torsion. Hence If $\gcd(\eta,\eta^{\prime})=1$ then $H_{i}(X^{3},\mathbb{Z})$ has no torsion for all $i$. ∎
Corollary 4.4.
For a prime $p$, there always exist a Plücker permutation $\sigma$ on $\{0,1,\dots,m\}$ such that $H_{i}(\sigma X^{3};\mathbb{Z})$ has no $p$-torsion for all $i$.
Theorem 4.5.
Proof.
For a prime $p$, choose a Plücker permutation $\sigma$ such that the minimum power of $p$ attains at $b_{\sigma(5)}$. Note that there are $4$ different Plücker permutation which satisfy the above condition. We can choose $\sigma$ such that $r^{\sigma(4)}\leq\{r^{\sigma(1)},r^{\sigma(2)},r^{\sigma(3)},r^{\sigma(4)}\}$. From the cofibration $L^{{}^{\prime}}(b_{\sigma(5)};\sigma{\bf{b}}^{(5)})\to\sigma X^{4}\hookrightarrow\mbox{Gr}_{\bf{b}}(2,4)$ we have
$\to H_{i}(\sigma X^{4})\to H_{i}(\mbox{Gr}_{\bf{b}}(2,4))\to\widetilde{H}_{i-1}(L^{\prime}(b_{\sigma(5)};\sigma{\bf{b}}^{(5)}))\to H_{i-1}(\sigma X^{4})\to.$
For $i=8$, we have
$\to 0\to H_{8}(\mbox{Gr}_{\bf{b}}(2,4))\to\mathbb{Z}\to 0\to.$
For $2\leq i\leq 7$, $\widetilde{H}_{i-1}(L^{\prime}(b_{\sigma(5)};\sigma{\bf{b}}^{(5)}))$ has no $p$-torsion. Thus $H_{i}(\mbox{Gr}_{\bf{b}}(2,4))$ has no $p$-torsion if the same is true for $H_{i}(\sigma X^{4})$. Similarly from the cofibration $L^{{}^{\prime}}(b_{\sigma(4)},\sigma{\bf{b}}^{(4)})\to\sigma X^{3}\hookrightarrow\sigma X^{4}$ we have
$\to\widetilde{H}_{i}(L^{\prime}(b_{\sigma(4)};\sigma{\bf{b}}^{4}))\to H_{i}(\sigma X^{3})\to H_{i}(\sigma X^{4})\to\widetilde{H}_{i-1}(L^{\prime}(b_{\sigma(4)};\sigma{\bf{b}}^{4}))\to.$
For $i=6$ we
$0\to H_{6}(\sigma X^{4})\to\mathbb{Z}\to 0$
Thus $H_{6}(\sigma X^{4})=\mathbb{Z}$. For $2\leq i\leq 5$, $H_{i}(\sigma X^{4})$ has no $p$-torsion if the same is true for $H_{i}(\sigma X^{3})$ as $\widetilde{H}_{i-1}(L^{\prime}(b_{\sigma(4)};\sigma{\bf{b}}^{(4)}))$ has no $p$-torsion. Using Lemma 4.3, $H_{i}(\sigma X^{3})$ has no torsion if $i\neq 2$. Thus $H_{i}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no torsion if $i\neq 2$. Hence the result follows using the universal coefficient Theorem for cohomology. ∎
Remark 4.6.
If $H_{2}(\sigma X^{3};\mathbb{Z})$ has no $p$-torsion for the same $\sigma$ that requires in the proof of Theorem 4.5 then $H_{i}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no $p$-torsion for every $i$.
Now we show that the theorem 4.1 is applicable to a huge class of weighted Grassmann orbifolds.
@@
    Let $\alpha,\beta,\gamma$ be three positive integers. Consider
    ${\bf{b}}=(\alpha\beta\gamma,\alpha\beta\gamma,(\alpha\beta-\alpha+1)\gamma,\alpha\gamma,\gamma,\gamma)\in(\mathbb{Z}_{\geq 1})^{6}.$
    Then $b_{i}$ divides $b_{j}$ for all $\lambda^{j}\preceq\lambda^{i}$ for $i\geq 3$. Thus the integral cohomology of $\mbox{Gr}_{\bf{b}}(2,4)$ has no torsion using Theorem 4.1.
@@
    ${\bf{b}}=(\gamma,(1+\beta)st\alpha,(1+s\beta)t\alpha,st\alpha,t\alpha,\alpha)\in(\mathbb{Z}_{\geq 1})^{6}.$
    In addition, if $\alpha+\gamma=t\alpha(1+s+s\beta)$ then ${\bf{b}}$ is a Plücker weight vector. The Plücker weight vector ${\bf{b}}$ satisfies the assumption of Theorem 4.1. Thus the integral cohomology of the corresponding $\mbox{Gr}_{\bf{b}}(2,4)$ has no torsion using Theorem 4.1.
Theorem 4.8.
[…]
Proof.
Since ${\bf{b}}$ is a Plücker weight vector we have $b_{1}+b_{4}=b_{2}+b_{3}$. If $b_{1}=b_{2}$ then $b_{3}=b_{4}$. Thus ${\bf{b}}=(b_{0},b_{1},b_{1},b_{3},b_{3},b_{5})$. For a prime $p$ we can assume $p$-th content of $b_{3}$ divides $p$-th content of $b_{1}$. Then by Theorem 4.1, $H^{*}(X^{4};\mathbb{Z})$ has no torsion. Otherwise consider the Plücker permutation $\sigma$ defined by $\sigma(1)=4$, $\sigma(4)=1$ and $\sigma(i)=i$ for $i\neq 1,4$. Then $\sigma{\bf{b}}=(b_{0},b_{3},b_{3},b_{1},b_{1},b_{5})$. Consider the weighted Grassmann orbifold $\mbox{Gr}_{\sigma{\bf{b}}}(2,4)$ to ensure that $H^{*}(X^{4};\mathbb{Z})$ has no torsion. Using the cohomology exact sequence in (4.1), $H^{*}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has maximum $b_{5}$ torsion. Consider the Plücker permutation $\sigma^{\prime}$ defined by $\sigma^{\prime}(0)=5$, $\sigma^{\prime}(5)=0$ and $\sigma^{\prime}(i)=i$ for $i\neq 0,5$. Then $\sigma^{\prime}{\bf{b}}=(b_{5},b_{1},b_{1},b_{3},b_{3},b_{0})$. By a similar argument $H^{*}(\mbox{Gr}_{\sigma^{\prime}{\bf{b}}}(2,4);\mathbb{Z})$ has maximum $b_{0}$ torsion. Since $(b_{0},b_{5})=1$ then $H^{*}(\mbox{Gr}_{\bf{b}}(2,4);\mathbb{Z})$ has no torsion. ∎
In the following example we discuss some classes of $\mbox{Gr}_{\bf{b}}(2,5)$ such that $H^{*}(\mbox{Gr}_{\bf{b}}(2,5);\mathbb{Z})$ has no torsion.

### 09 Mar 2026 v3

@@ 3. Classification of weighted Grassmann orbifolds upto certain homeomorphism
Two weighted projective spaces $\mathbb{W}P(b_{0},\dots,b_{m})$ and $\mathbb{W}P(c_{0},\dots,c_{m})$ are homeomorphic if and only if $(b_{0},\dots,b_{m})$ and $(c_{0},\dots,c_{m})$ have same normalized form up to an order.
Lemma 3.10.
Let $k\geq 2$ and ${\bf b}=(b_{0},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ be a primitive Plücker weight vector such that $m+1={n\choose k}$. Then $(b_{0},\dots,b_{m})$ is a normalized Plücker weight vector.
Proof.
Let $p$ be a prime and ${\bf b}=(b_{0},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ be a primitive Plücker weight vector. Then there exist at least one $b_{i}$ such that $p$ does not divide $b_{i}$. Let $\lambda^{i}=(\lambda_{1}^{i},\lambda_{2}^{i},\lambda_{3}^{i},\dots,\lambda_{k}^{i})$. We consider $1\leq q,r\leq n$ such that none of $q$ and $r$ is in $\lambda^{i}$. Define $\lambda^{j}=(q,r,\lambda_{3}^{i},\dots,\lambda_{k}^{i})$, $\lambda^{s}=(q,\lambda_{2}^{i},\lambda_{3}^{i},\dots,\lambda_{k}^{i})$, $\lambda^{\ell}=(\lambda_{1}^{i},r,\lambda_{3}^{i},\dots,\lambda_{k}^{i})$. Then using Remark 2.4, $b_{i}+b_{j}=b_{s}+b_{\ell}$. Since $p$ does not divide $b_{i}$, $p$ does not divide at least one of $b_{j},b_{k},b_{\ell}$. This is true for all prime $p$. Hence the proof. ∎
Define
@@
Then cardinality of $Ad(\lambda^{i})$ is k(nk)k(n-k). Corresponding to $Ad(\lambda^{i})$ we define
${\bf b}_{(i)}:=\{{b_{j}}:\lambda^{j}\in Ad(\lambda^{i})\}\text{ and }\overline{\bf b}_{(i)}:={\bf b}_{(i)}\cup\{b_{i}\}.$
For $i\in\{0,1,\dots,m\}$ define
(3.2) $U_{i}:=\{[z_{0},\dots,z_{m}]\in\mbox{Gr}_{\bf b}(k,n)~|~z_{i}\neq 0\}.$
From Section 2, it follows that $U_{i}\cong\mathbb{C}^{k(n-k)}/G(b_{i})$. Note that $U_{i}$ is an open cone over $L^{\prime}(b_{i};{\bf b}_{(i)})$. Let t:=k(nk)t:=k(n-k). Using [15, Theorem 2] we have
(3.3) $H^{q}(L^{\prime}(b_{i};{\bf{b}}_{(i)});\mathbb{Z})=\begin{cases}\mathbb{Z}&\quad\text{for}~q=0,2t-1\\ \mathbb{Z}_{\mu}&\quad\mu={l_{e}{\overline{\bf{b}}_{(i)}}}/{l_{e}{{\bf{b}}_{(i)}}}~\text{if}~q=2e~\text{for}~1\leq e\leq t-1\\ \{0\}&\quad\text{ otherwise}.\end{cases}$
@@
(3.4) $l_{e}{\bf{b}}_{(i)}=\prod_{p:~\text{prime}}\prod_{i=1}^{e}p^{r_{t-i+1}}~\text{for}~1\leq e\leq t-1.$
@@
Let $k\geq 2$ and ${\bf b}=(b_{0},\dots,b_{m})\in(\mathbb{Z}_{\geq 1})^{m+1}$ be a primitive Plücker weight vector such that $m+1={n\choose k}$. Then $\overline{\bf{b}}_{(i)}$’s are primitive vectors for all $i=0,1,\dots,m$.
@@
Fix one $i\in\{0,1,\dots,m\}$. If $b_{j}\notin\overline{\bf b}_{(i)}$ then $|\lambda^{j}\cap\lambda^{i}|=k-u\leq k-2$. Without loss of generality assume that $\lambda^{i}=(\lambda^{i}_{1},\dots,\lambda^{i}_{u},\lambda^{i}_{u+1},\dots,\lambda^{i}_{k})$ and $\lambda^{j}=(\lambda^{j}_{1},\dots,\lambda^{j}_{u},\lambda^{i}_{u+1},\dots,\lambda^{i}_{k})$, where $\lambda^{i}_{1}\neq\lambda^{j}_{1},$ $\dots,$ $\lambda^{i}_{u}\neq\lambda^{j}_{u}$. Consider Schubert symbols $\lambda^{s_{1}}=(\lambda^{j}_{1},\lambda^{i}_{2}\dots,\lambda^{i}_{u},\lambda^{i}_{u+1},\dots,\lambda^{i}_{k}),$ $\dots,\lambda^{s_{u}}=(\lambda^{i}_{1},\dots,\lambda^{i}_{u-1},\lambda^{j}_{u},\lambda^{i}_{u+1},\dots,\lambda^{i}_{k})$. Then $|\lambda^{i}\cap\lambda^{s_{v}}|=k-1$ and $b_{s_{v}}\in{\bf b}_{(i)}$ for all $v=1,2,\dots,u$. Note that
@@
If $d$ divides $b_{i}$ and every element of ${\bf b}_{(i)}$ then $d$ also divides $b_{j}$. Hence, $\overline{\bf b}_{(i)}$ is not a primitive vector contradicts the fact that $\bf b$} is a primitive Plücker weight vector. ∎
We recall the coordinate elements $e_{\lambda^{i}}]\in U_{i}$, for $i=0,1,\dots,m$, where $e_{\lambda^{i}}$ is defined in Subsection 2.1.
@@
Let $k\geq 2$ and ${\bf b}=(b_{0},\dots,b_{m})$ be a primitive Plücker weight vector. Then
@@
We recall $U_{i}$ from (3.2). Using excision theorem
@@
Following the notation of (3.3), we need to prove
@@
Let $k\geq 2$. If there exists a homeomorphism $f$ between two weighted Grassmann orbifolds ${\rm{Gr}}_{\bf b}(k,n)$ and ${\rm{Gr}}_{\bf c}(k,n)$ such that $f$ maps every coordinate element $e_{\lambda^{i}}]$ to some coordinate element $e_{\lambda^{j}}]$, where $i,j\in\{0,1,\dots,m\}$, then $\bf b$} and $\bf c$} are same up to a scalar multiplication and a permutation.
@@
We recall the $(\mathbb{C}^{*})^{n}$-action on $\mbox{Gr}_{\bf b}(k,n)$ from (3.1) and note that the fixed point set of $\mbox{Gr}_{\bf b}(k,n)$ with respect to the $(\mathbb{C}^{*})^{n}$-action is $\{[e_{\lambda^{i}}]\}_{i=0}^{m}$.
@@
Let $k\geq 2$ and $G$ be a group acts on a weighted Grassmann orbifold $\mbox{Gr}_{\bf b}(k,n)$ with the fixed point set $\{[e_{\lambda^{i}}]\}_{i=0}^{m}$. If ${\rm{Gr}}_{\bf b}(k,n)$ is weakly $G$-equivariantly homeomorphic to ${\rm{Gr}}_{\bf c}(k,n)$, then $\bf b$ and $\bf c$ are the same upto a scalar multiplication and a permutation.
@@
In the view of Lemma 3.10, Lemma 3.11 and Lemma 3.12 we propose the following conjecture.
Conjecture 3.15.
Let $k\geq 2$. If two weighted Grassmann orbifolds ${\rm{Gr}}_{\bf b}(k,n)$ and ${\rm{Gr}}_{\bf c}(k,n)$ are homeomorphic then then $\bf b$} and $\bf c$} are same up to a scalar multiplication and a Plücker permutation.
@@ 4. Torsion’s in the integral cohomology of weighted Grassmann orbifold
In this section, we prove that for a prime $p$ the integral cohomologies of weighted Grassmann orbifolds have no $p$-torsion under some assumptions. We discuss some non-trivial classes of weighted Grassmann orbifolds whose integral cohomologies have no torsions. In particular, the integral cohomology of a divisive weighted Grassmann orbifold is torsion-free and concentrated in even degrees. Recall $I(\lambda^{i})$ from (2.15).
@@
Let $p$ be a prime number. If for every $j\leq m-3$, the $p$-th content of $b_{j}$ divides $d^{\prime}(\lambda^{j})-1$ many elements of the set $\{b_{\ell}:\lambda^{\ell}\in I(\lambda^{j})\}$ then the integral cohomology of ${\rm{Gr}}_{\bf b}(k,n)$ has no $p$-torsion.
@@
Recall the building sequence
$X^{0}\subset X^{1}\subset\dots\subset X^{m}=\mbox{Gr}_{\bf b}(k,n),$
induced from the $q$-CW complex structure as in (2.14). We prove that the integral cohomology of $X^{i}$ has no torsion for all $0\leq i\leq m$ by using induction on $i$. Note that $X^{0}=\{pt\},X_{1}=\mathbb{W}P(b_{m-1},b_{m})$ and $X^{2}=\mathbb{W}P(b_{m-2},b_{m-1},b_{m})$ are weighted projective spaces. Thus, the integral cohomology of $X^{i}$ has no torsion for all $0\leq i\leq 2$ by [15, Theorem 1]. Let us assume $H^{*}(X^{i-1};\mathbb{Z})$ has no $p$-torsion. We recall the following cofibration from Proposition 2.12
$L^{\prime}(b_{i};{\bf{b}}^{(i)})\hookrightarrow X^{i-1}\to X^{i},$
where ${\bf b}^{(i)}$ is defined prior to Proposition 2.12. This will induce the following exact sequence in cohomology
(4.1) $\to\widetilde{H}^{q-1}(L^{\prime}(b_{i};{\bf{b}}^{(i)});\mathbb{Z})\to H^{q}(X^{i};\mathbb{Z})\to H^{q}(X^{i-1};\mathbb{Z})\to\widetilde{H}^{q}(L^{\prime}(b_{i};{\bf{b}}^{(i)});\mathbb{Z})\to.$
Using [15, Theorem 2] we have
$H^{q}(L^{\prime}(b_{i};{\bf{b}}^{(i)});\mathbb{Z})=\begin{cases}\mathbb{Z}&\quad\text{for}~q=0,2d^{\prime}(\lambda^{i})-1\\ \mathbb{Z}_{\mu}&\quad\mu={l_{e}{\overline{\bf{b}}^{(i)}}}/{l_{e}{{\bf{b}}^{(i)}}}~\text{if}~q=2e~\text{for}~1\leq e\leq d^{\prime}(\lambda^{i})-1\\ \{0\}&\quad\text{ otherwise},\end{cases}$
where ${\overline{\bf{b}}^{(i)}}:={\bf{b}}^{(i)}\cup\{b_{i}\}.$ For any prime $p$ the $p$-th content in $l_{e}{{\bf{b}}^{(i)}}$ is the product of $e$ many maximum $p$-th contents of the elements in set $\{b_{\ell}\colon\lambda^{\ell}\in I(\lambda^{i})\}$ as in (3.4). Thus the $p$-th content of $\mu$ is 1, if the $p$-th content of $b_{i}$ divides $e$ many element of the set $\{b_{\ell}:\lambda^{\ell}\in I(\lambda^{i})\}$. Now the maximum value of $e$ can be $d^{\prime}(\lambda^{i})-1$. Thus by our assumption $H^{*}(L^{\prime}(b_{i};{\bf{b}}^{(i)});\mathbb{Z})$ has no $p$-torsion. Thus $H^{*}(X^{i},\mathbb{Z})$ has no $p$-torsion using (4.1) and induction hypothesis. Hence, by induction the integral cohomology of $X^{m}=\mbox{Gr}_{\bf b}(k,n)$ has no $p$-torsion. ∎
Corollary 4.2.
For a prime $p$, if there exists a Plücker permutation $\sigma$ on $\{0,\dots,m\}$ such that for all $j\leq m-3$, the $p$-th content of $b_{\sigma(j)}$ divides $d^{\prime}(\lambda^{j})-1$ many elements of the set $\{b_{\sigma(\ell)}:\lambda^{\ell}\in I(\lambda^{j})\}$ then the integral cohomology of ${\rm{Gr}}_{\bf b}(k,n)$ has no $p$-torsion.
Proof.
Let $\sigma$ be a a Plücker permutation. Consider the building sequence of $\mbox{Gr}_{\bf b}(k,n)$ induced from the $q$-CW complex structure following Remark 3.5. Now the proof followes using the same argument as in the Proof of Theeorem 4.1. ∎
@@
    Let $\alpha,\beta,\gamma$ be three positive integers. Consider a Plücker vector
    ${\bf{b}}=(\gamma,\gamma,\alpha\gamma,(\alpha\beta-\alpha+1)\gamma,\alpha\beta\gamma,\alpha\beta\gamma)\in(\mathbb{Z}_{\geq 1})^{6}.$
    Then $b_{i}$ divides $b_{j}$ for all $\lambda^{j}\succeq\lambda^{i}$ for $i\leq 2$. Thus the integral cohomology of the $\mbox{Gr}_{\bf{b}}(2,4)$ has no torsion using Theorem 4.1.
@@
    ${\bf{b}}=(\alpha,t\alpha,st\alpha,(1+s\beta)t\alpha,(1+\beta)st\alpha,\gamma)\in(\mathbb{Z}_{\geq 1})^{6}.$
    In addition, if $\alpha+\gamma=t\alpha(1+s+s\beta)$ then $\bf{b}$} is a Plücker weight vector. The integral cohomology of the $\mbox{Gr}_{\bf{b}}(2,4)$ has no torsion using Theorem 4.1.